Mazeppa Corporation sells relays at a selling price of $28 per unit. The company's cost per unit, based on full capacity of 160,000 units, is as follows:

Direct materials $ 6
Direct labor 4
Overhead (2/3 of which is variable) 9

Mazeppa has been approached by a distributor in Montana offering to buy a special order consisting of 30,000 relays. Mazeppa has the capacity to fill the order. However, it will incur an additional shipping cost of $2 for each relay it sells to the distributor.

a-1. Assume that Mazeppa is currently operating at a level of 100,000 units. Show the calculation for the unit price to charge the distributor which will generate an increase in operating income of $2 per unit?

a-2. What is your interpretation of the changes to the contribution margin per unit and the operating income on account of the unit price charged to the distributor?

b-1. Assume that Mazeppa is currently operating at full capacity. Show the calculation for the unit price to charge the distributor which will generate an increase in operating income of $60,000 more than it would be without accepting the special order?

b-2. What is your interpretation of the changes to the contribution margin per unit and the operating income on account of the unit price charged to the distributor?
Answer #1
a1) In the position of unused capacity, the fixed overheads are
not applicable on special order. Only variable costs are considered.
Thus, the unit price to charge the distributor for 30000 relays is:
Variable costs + expected operating income per unit =
(Direct Mat. + Direct Lab. + Variable overhead + Shipping cost) +
expected operating income per unit
The unit price = (6+4+6+2) + 2 = 18 + 2 =$20 per unit
a2) The contribution margin per unit = Unit price - Unit variable costs =
$20 - (6+4+6+2) = $2 per unit and operating income per unit is $2.
Normal contribution per unit is $28 - $16 = $12 per unit and operating
income per unit is $28 - total costs $19 = $9.
b1) As Mazeppa is operating at full capacity, so the price charged to distributor
is calculated as :
The operating income without special order = (28 - 19)* 160000 = $1440000
The desired operating income = 1440000 + 60000 = 1500000
The installed units sold except special order = 160000 - 30000 = 130000
Normal operating income = 130000 * $9 = 1170000
Operating income expected of special order = Desired operating income - Normal operating income =
1500000 - 1170000 = 330000
Per unit expected operating income of special order = 330000 / 30000 = $11 per unit
Variable costs per unit of special order = $18
Required special order price per unit (total units 30000) = $11 + $18 = $29
b2) The contribution margin per unit = Unit price - Unit variable costs =
$29 - $18 = $11 per unit and operating income per unit is $11.
Normal contribution per unit is $28 - $16 = $12 per unit and operating
income per unit is $28 - total costs $19 = $9.